OpenStudy (anonymous):

solve 3 + sqrt z-7=sqrt z+8

OpenStudy (anonymous):

is that right ? check again!

OpenStudy (anonymous):

I don't think your question is right.

jhonyy9 (jhonyy9):

3+sqrt(z-7)=sqrt(z+8) squared 9+6sqrt(z-7)+7=z+8 6sqrt(z-7)=z-8 squared 36(z-7)=z2 -16z +64 36z - 252=z2 -16z +64 z2 -52z +316=0 z1,2=(52+/- sqrt(2704 - 1264))/2 =(52+/- sqrt1440)/2 =(52+/- 12sqrt10)/2 = = 26+/- 6sqrt10

OpenStudy (anonymous):

can you please right the equation as it is because it looks vague